Sina+cosb=1/√2 & cosa+sinb=√2/√3 then find the value of : (1) sin(a+b) (2) tan(a-b/2)

Grade:11

1 Answers

Vikas TU
7 years ago
Sina+cosb=1/√2
Sina = 1/√2 – cosb
squaring both sides,
(sina)^2  =  1/2 + (cosb)^2 – 2cosb/√2
1 – (cosa)^2 = 1/2 + (cosb)^2 – 2cosb/√2
1/2 – (cosa)^2 = (cosb)^2 – 2cosb/√2.................(1) 
This is the eqn. 1
Do it same squaring method in  cosa+sinb=√2/√3
and bring the final eqn. in cosa and cosb terms
then after solve those eqns. andd get cosa and cosb values.